OpenStudy (anonymous):

i don't understand how to do this, so please show steps. Factor the polynomial. 64x^(3)-y^(6). Is this difference of cubes?

OpenStudy (anonymous):

If not possible thn say 'NF'

OpenStudy (mertsj):

\[64x^3-y^6=[(4x)^3-(y^2)^3]\]

OpenStudy (mertsj):

So we see that it is the difference of two cubes. Use: \[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (mertsj):

And replace each a with 4x and each b with y^2

OpenStudy (mertsj):